OpenStudy (anonymous):

How do you put -3i+7 / 5 in a+bi form?

OpenStudy (anonymous):

Well, you can separate the numerator like this: \[\dfrac {-3} {5} i+\dfrac {7} {5}\]

OpenStudy (anonymous):

it has to be in a+bi form.

OpenStudy (anonymous):

well, 7/5 plus -3/5 i, where a= 7/5 and b=-3/5, a and b just got to be real.